# The lattice packing problem in dimension 9 by Voronoi’s algorithm

### v1

@@ 6.2.3. Extra symmetries and double coset decomposition
After the enumeration of this work was completed, we finally implemented ascending chains of subgroups and double coset enumeration in Permutalib [DS24] inspired by the implementation of GAP [GAP24]. This turns out to be in most cases the most efficient technique.
@@ 6.2.4. Saving Bank
Suppose that we compute the dual description of facets $F_{1}$ and $F_{2}$ by using the adjacency decomposition method and that $F_{1}$ and $F_{2}$ share a common ridge $R$. Then the dual description of $R$ could be computed two times which is a waste of resources. In case the algorithm recurses the dual description of lower-dimensional faces this could even increase more. Furthermore, some of the ridges $R$ could simply be equivalent as a polyhedral cone and thus have equivalent dual descriptions. Therefore, we have a memoization process which stores the computed dual descriptions in a saving bank if they were expensive to compute, following the approach from [DSSV07]. To quickly determine if an equivalent dual description has been stored already, up to restricted linear equivalence, we use the canonical form as detailed in section 5.3. If the dual description under symmetry has been stored with a different symmetry group we use the double coset decomposition techniques treated in section 6.2.3.
@@ 6.3. Implementation details and improvements
The algorithmic improvements alone were not sufficient, so we will now discuss some of the implementation details and improvements. Ignoring the improvements to parallelism, our implementation effort along with the algorithmic improvements reduced the dual description computation for $\operatorname{Vor}(\mathsf{E}_{8})$ from several months, as reported in [DSSV07], down to only $9$ hours on a single core. Additionally, our highly parallel implementation made the much larger cases brought up by Voronoi’s algorithm in dimension $9$ feasible, on which we will report more in section 7.
@@ 6.3.1. C++ implementation
For the classification of the perfect forms of dimension $8$ the main bottleneck was the dual description of $\operatorname{Vor}(\mathsf{E}_{8})$ which was performed using an implementation of Recursive ADM in the Polyhedral package [DS15] in the GAP system for computational discrete algebra [GAP24]. While GAP has very efficient algorithms for permutation groups it is not sufficiently fast and capable for the large scale of our computation. Therefore, as part of this work the Recursive ADM algorithm, along with all necessary permutation group features of GAP, have been ported to a C++ implementation. This implementation has been developed publicly at [DSvW25b]. It depends on the Eigen, GMP and Boost [GJ+10, Gt23, The25] libraries for handling matrices, exact arithmetic, and many miscellaneous functions such as serialization respectively. Furthermore, it depends on Nauty [MP22] for the computation of canonical graphs, and uses PPL [BHZ08] or LRS [Avi15] for basic dual description computations.
@@
When computing the ridges $R_{i}$ with a direct dual description method that does not consider any symmetries, such as with PPL [BHZ08] or LRS [Avi15], one often directly obtains both the extreme rays included in $R_{i}$, and additionally the face-defining inequality $f_{i}$ of $R_{i}$. The flipping process, which uses both these data, can then immediately be applied.
@@
We noticed however for our cases that the face-defining inequalities can typically be given by integer vectors with relatively small entries. Therefore, we computed a kernel vector $a\in\mathbb{F}_{p}^{n}$ modulo some large prime $p\approx 2^{31}$, which is much faster as operations over this field can be implemented efficiently using $64$-bit integer arithmetic. Then, using rational lifting, via continuous fraction methods [WGD82] or Lagrange reduction, we lift each vector coefficient $a_{i}$ in $\mathbb{F}_{p}$ to an appropriate rational $\frac{b_{i}}{c_{i}}\in\mathbb{Q}$ such that $a_{i}=b_{i}c_{i}^{-1}\bmod{p}$ and such that $b_{i}$ and $c_{i}$ are small ($|b_{i}|,|c_{i}|<\sqrt{p}$). We then remove the common denominator to obtain an integer vector. Typically this vector lies in the kernel which we check. In case the check fails we fall back to the original computation over the rationals, however we have not observed that this was required. The resulting conversion technique leads to a $10\times$ or larger speed-up compared to the original approach.
@@ 6.3.3. Parallel adjacency method via MPI
Just as for our implementation of Voronoi’s algorithm we used MPI [MPI25] to scale the algorithm up to hundreds of processors. On a high level our approach is also similar to our approach there so we will be brief about it. We mostly parallelized the Recursive ADM algorithm on the top level, i.e., the facet orbits are distributed over the processors and treated separately. The canonical function to represent the facets detailed in  section 5.4 is, along with a hash function, used to send each facet to the appropriate processors, just as done with the canonical perfect forms.
@@ 7. Results
In this section we will consider the cost of our computations, some data resulting from our enumeration, and the corollaries resulting from those. Many of the questions we try to answer are inspired by a collection of open problems posed by Martinet [Mar15]. We will generally denote forms by $Q_{a[,b]}$, where $a$ indicates half the number of minimum vectors, and $b$ optionally indicates the specific lattice it is related to or the size of its automorphism group.
@@ 7.1. Classification of perfect forms in dimension 9
We completely finished the execution of Voronoi’s algorithm and thereby found a complete list of 2 237 251 040 non-similar perfect forms in dimension $9$, leading to theorem 1. Among this complete list of perfect lattices the laminated lattice $\mathsf{\Lambda}_{9}$ [Cha46, CS82] is the densest lattice. By Voronoi’s theory, in particular lemma 2, we can conclude from this in corollary 1 that $\mathsf{\Lambda}_{9}$ is the unique densest lattice packing in dimension $9$ and we have the Hermite constant $\gamma_{9}=2$.
We now consider some further properties of the classification. In table 3 we show the number of non-similar forms by their kissing number $|\operatorname{Min}(Q)|$. From $|\operatorname{Min}(Q)|\geq 2\cdot 63$ this matches the partial enumeration from [vW18, Table 1], so no new high-incidence forms were found. This is expected as the high-incidence forms are often highly connected and found early. Just as in dimension $8$ the minimal kissing number $n=\frac{1}{2}d(d+1)$ for a $d$-dimensional perfect form is attained by most, more than half, of the perfect forms. Above this we roughly see an exponential decrease as the kissing number increases. As was already known by Watson [Wat71], the maximum kissing number in dimension $9$ is $2\cdot 136$ attained by $\mathsf{\Lambda}_{9}$.
In table 6 in the Appendix we display the number of non-similar forms by the order $|\operatorname{Aut}(Q)|$ of their automorphism group. More than $96.7\%$ of them have a trivial automorphism group of order $2$, and more than $99.9\%$ of them have an order of at most $4$. The largest automorphism group has order $2^{9}\cdot 9!$ and is attained by the root lattice $\mathsf{D}_{9}$ [CS13, p. 117].
@@ 7.2. The computational cost and high-incidence cases
Surprisingly, the highest-incidence case with incidence $136$ corresponding to the Laminated lattice $\mathsf{\Lambda}_{9}$, was not the most computationally intensive. We see in table 4 that this is mainly the case due to the many extra linear automorphisms of the cone, i.e., its automorphism group is a factor $128$ larger than the subgroup induced by $\operatorname{Aut}(Q_{136,\mathsf{\Lambda}_{9}})$, as was already observed in [vW18]. Furthermore, the perfect form with incidence $129$ does not even appear in this list as its dual description easily follows from that of $\mathsf{E}_{8}$, again see [vW18].
@@ 7.3. Perfect forms only connected via high-incidence cases
$Q_{74,256}$$Q_{45,1440}$$Q_{90,\mathsf{\kappa}_{9}}$$Q_{45,7680}$$Q_{45,1536}$$Q_{45,2304}$$Q_{45,11520}$$Q_{45,161280}$$Q_{45,\mathsf{A}_{9}^{5}}$$Q_{136,\mathsf{\Lambda}_{9}}$$Q_{45,\mathsf{A}_{9}^{2}}$
@@
We consider all forms that were not found yet after incrementally exploring all forms until a kissing number of $2\cdot 60$. There are only three perfect forms with a large kissing number that lead to new perfect forms, namely forms that we denote by $Q_{74,256},Q_{90,\mathsf{\kappa}_{9}}$ and $Q_{136,\mathsf{\Lambda}_{9}}$ with a kissing number of respectively $2\cdot 74,2\cdot 90$ and $2\cdot 136$. The latter two correspond to the lattices $\mathsf{\kappa}_{9}$ [PP93] and the laminated lattice $\mathsf{\Lambda}_{9}$ [Cha46, CS82] respectively, both sections of the Leech lattice $\Lambda_{24}$ [Lee67].
The new perfect forms and their connections in Voronoi’s graph are presented in fig. 4. Being only connected via these high-incidence perfect forms already makes them interesting, but in fact quite some of these lattices have other special properties. All these new perfect forms have the minimal possible kissing number of $2\cdot 45$, but all have a large automorphism group which is typically not the case for such low-incidence forms. The perfect form $Q_{45,\mathsf{A}_{9}^{2}}$ is the unique perfect form that is only connected to the perfect lattice $\mathsf{\Lambda}_{9}$, and corresponds to Coxeter’s lattice $\mathsf{A}_{9}^{2}$ [Cox51]. We will see later that this lattice is the only $9$-dimensional perfect lattice whose minimal vectors do not span the lattice. Then there is the perfect form $Q_{45,\mathsf{A}_{9}^{5}}$ [Cox51, MB04], dual to $Q_{45,\mathsf{A}_{9}^{2}}$, which is only connected in the Voronoi graph via two and three steps with $Q_{90,\mathsf{\kappa}_{9}}$ and $Q_{136,\mathsf{\Lambda}_{9}}$ respectively. We will show later that this lattice is in fact the only so-called hollow perfect lattice in dimension $9$. The perfect forms $Q_{45,\mathsf{A}_{9}^{5}},Q_{45,\mathsf{A}_{9}^{2}}$ and $Q_{45,7680}$ all satisfy $\gamma^{\prime}(Q)^{2}=\lambda_{1}(Q)\cdot\lambda_{1}(Q^{-1})=\frac{16}{5}$, the maximum known value for the Bergé-Martinet invariant $\gamma^{\prime}(Q)$ in dimension $9$ [BM89]. Along with the perfect form $Q_{45,1440}$, these $4$ forms are both extreme and dual-extreme, i.e., they achieve a local-maximum for the Bergé-Martinet invariant. Furthermore, the perfect form $Q_{45,161280}$ achieves a rather high value of $\gamma^{\prime}(Q_{45,161280})=37/12$, but is not locally-optimal, due to its better neighbour $Q_{45,\mathsf{A}_{9}^{5}}$.
@@ 7.4. Possible kissing numbers in dimension 99
By definition $Q$ lies on the face of the Rhyskov polyhedra $\mathcal{P}_{\lambda_{1}(Q)}^{9}$ defined by the facets corresponding to $\operatorname{Min}(Q)/\{\pm\}$. The incidence of the face is $\frac{1}{2}|\operatorname{Min}(Q)|$ and thus corresponds to half the kissing number of $Q$. This face is either a vertex, i.e., $Q$ is a perfect form, or the face contains a vertex corresponding to a perfect form $Q^{\prime}$. In the first case we know by our full classification that $|\operatorname{Min}(Q)|\in\{45,\ldots,82,84,90,91,99,129,136\}$. Furthermore, we know that $1,\ldots,44$ are attained in dimension $8$ [DSSV07] which by a direct-sum construction can be lifted to dimension $9$. In the second case the face can be identified with a non-trivial face of the tangent cone $\mathcal{P}(Q^{\prime})$ at $Q^{\prime}$.
@@
In [vW18] it was shown that $\operatorname{Vor}(Q_{129})$ can be decomposed into the Voronoi domain $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ of the $8$-dimensional root lattice $\mathsf{E}_{8}$ with $120$ extreme rays and a linearly independent simplicial cone with $9$ extreme rays. As a result, all faces of $\operatorname{Vor}(Q_{129})$ can be described as a sum of a face of $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ and a face of the simplicial cone. If we consider all faces containing $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ we thus obtain non-trivial faces of incidence $120+k$ for $k=0,\ldots,8$. The largest non-trivial face of $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ has incidence $75$, therefore all other faces of $\operatorname{Vor}(Q_{129})$ not containing $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ have incidence at most $75+9=84$.

### v2

@@ 6.2.3. Extra symmetries and double coset decomposition
After the enumeration of this work was completed, we finally implemented ascending chains of subgroups and double coset enumeration in Permutalib [permutalib] inspired by the implementation of GAP [GAP]. This turns out to be in most cases the most efficient technique.
@@ 6.2.4. Saving Bank
Suppose that we compute the dual description of facets $F_{1}$ and $F_{2}$ by using the adjacency decomposition method and that $F_{1}$ and $F_{2}$ share a common ridge $R$. Then the dual description of $R$ could be computed two times which is a waste of resources. In case the algorithm recurses the dual description of lower-dimensional faces this could even increase more. Furthermore, some of the ridges $R$ could simply be equivalent as a polyhedral cone and thus have equivalent dual descriptions. Therefore, we have a memoization process which stores the computed dual descriptions in a saving bank if they were expensive to compute, following the approach from [PerfectDim8]. To quickly determine if an equivalent dual description has been stored already, up to restricted linear equivalence, we use the canonical form as detailed in section 5.3. If the dual description under symmetry has been stored with a different symmetry group we use the double coset decomposition techniques treated in section 6.2.3.
@@ 6.3. Implementation details and improvements
The algorithmic improvements alone were not sufficient, so we will now discuss some of the implementation details and improvements. Ignoring the improvements to parallelism, our implementation effort along with the algorithmic improvements reduced the dual description computation for $\operatorname{Vor}(\mathsf{E}_{8})$ from several months, as reported in [PerfectDim8], down to only $9$ hours on a single core. Additionally, our highly parallel implementation made the much larger cases brought up by Voronoi’s algorithm in dimension $9$ feasible, on which we will report more in section 7.
@@ 6.3.1. C++ implementation
For the classification of the perfect forms of dimension $8$ the main bottleneck was the dual description of $\operatorname{Vor}(\mathsf{E}_{8})$ which was performed using an implementation of Recursive ADM in the Polyhedral package [Polyhedral] in the GAP system for computational discrete algebra [GAP]. While GAP has very efficient algorithms for permutation groups it is not sufficiently fast and capable for the large scale of our computation. Therefore, as part of this work the Recursive ADM algorithm, along with all necessary permutation group features of GAP, have been ported to a C++ implementation. This implementation has been developed publicly at [PolyhedralCpp]. It depends on the Eigen, GMP and Boost [eigenweb, GMP3, boostweb] libraries for handling matrices, exact arithmetic, and many miscellaneous functions such as serialization respectively. Furthermore, it depends on Nauty [nauty] for the computation of canonical graphs, and uses PPL [BagnaraHZ08SCP] or LRS [lrs] for basic dual description computations.
@@
When computing the ridges $R_{i}$ with a direct dual description method that does not consider any symmetries, such as with PPL [BagnaraHZ08SCP] or LRS [lrs], one often directly obtains both the extreme rays included in $R_{i}$, and additionally the face-defining inequality $f_{i}$ of $R_{i}$. The flipping process, which uses both these data, can then immediately be applied.
@@
We noticed however for our cases that the face-defining inequalities can typically be given by integer vectors with relatively small entries. Therefore, we computed a kernel vector $a\in\mathbb{F}_{p}^{n}$ modulo some large prime $p\approx 2^{31}$, which is much faster as operations over this field can be implemented efficiently using $64$-bit integer arithmetic. Then, using rational lifting, via continuous fraction methods [wang1982p] or Lagrange reduction, we lift each vector coefficient $a_{i}$ in $\mathbb{F}_{p}$ to an appropriate rational $\frac{b_{i}}{c_{i}}\in\mathbb{Q}$ such that $a_{i}=b_{i}c_{i}^{-1}\bmod{p}$ and such that $b_{i}$ and $c_{i}$ are small ($|b_{i}|,|c_{i}|<\sqrt{p}$). We then remove the common denominator to obtain an integer vector. Typically this vector lies in the kernel which we check. In case the check fails we fall back to the original computation over the rationals, however we have not observed that this was required. The resulting conversion technique leads to a $10\times$ or larger speed-up compared to the original approach.
@@ 6.3.3. Parallel adjacency method via MPI
Just as for our implementation of Voronoi’s algorithm we used MPI [MPIforum] to scale the algorithm up to hundreds of processors. On a high level our approach is also similar to our approach there so we will be brief about it. We mostly parallelized the Recursive ADM algorithm on the top level, i.e., the facet orbits are distributed over the processors and treated separately. The canonical function to represent the facets detailed in  section 5.4 is, along with a hash function, used to send each facet to the appropriate processors, just as done with the canonical perfect forms.
@@ 7. Results
In this section we will consider the cost of our computations, some data resulting from our enumeration, and the corollaries resulting from those. Many of the questions we try to answer are inspired by a collection of open problems posed by Martinet [Martinet2015]. We will generally denote forms by $Q_{a[,b]}$, where $a$ indicates half the number of minimum vectors, and $b$ optionally indicates the specific lattice it is related to or the size of its automorphism group.
@@ 7.1. Classification of perfect forms in dimension 9
We completely finished the execution of Voronoi’s algorithm and thereby found a complete classification of 2 237 251 040 non-similar perfect forms in dimension $9$, leading to theorem 1. Among this complete list of perfect lattices the laminated lattice $\mathsf{\Lambda}_{9}$ [chaundy1946arithmetic, conway1982laminated] is the densest lattice. By Voronoi’s theory, in particular lemma 2, we can conclude from this in corollary 1 that $\mathsf{\Lambda}_{9}$ is the unique densest lattice packing in dimension $9$ and we have the Hermite constant $\gamma_{9}=2$.
We now consider some further properties of the classification. In table 3 we show the number of non-similar forms by their kissing number $|\operatorname{Min}(Q)|$. From $|\operatorname{Min}(Q)|\geq 2\cdot 63$ this matches the partial enumeration from [vanWoerdenMasterThesis, Table 1], so no new high-incidence forms were found. This is expected as the high-incidence forms are often highly connected and found early. Just as in dimension $8$ the minimal kissing number $n=\frac{1}{2}d(d+1)$ for a $d$-dimensional perfect form is attained by most, more than half, of the perfect forms. Above this we roughly see an exponential decrease as the kissing number increases. As was already known by Watson [G1971], the maximum kissing number in dimension $9$ is $2\cdot 136$ attained by $\mathsf{\Lambda}_{9}$.
In table 6 in the Appendix we display the number of non-similar forms by the order $|\operatorname{Aut}(Q)|$ of their automorphism group. More than $96.7\%$ of them have a trivial automorphism group of order $2$, and more than $99.9\%$ of them have an order of at most $4$. The largest automorphism group has order $2^{9}\cdot 9!$ and is attained by the root lattice $\mathsf{D}_{9}$ [conway2013sphere, p. 117].
@@ 7.2. The computational cost and high-incidence cases
Surprisingly, the highest-incidence case with incidence $136$ corresponding to the Laminated lattice $\mathsf{\Lambda}_{9}$, was not the most computationally intensive. We see in table 4 that this is mainly the case due to the many extra linear automorphisms of the cone, i.e., its automorphism group is a factor $128$ larger than the subgroup induced by $\operatorname{Aut}(Q_{136,\mathsf{\Lambda}_{9}})$, as was already observed in [vanWoerdenMasterThesis]. Furthermore, the perfect form with incidence $129$ does not even appear in this list as its dual description easily follows from that of $\mathsf{E}_{8}$, again see [vanWoerdenMasterThesis].
@@ 7.3. Perfect forms only connected via high-incidence cases
$Q_{72,D_{9}}$$Q_{45,A_{9}}$$Q_{74,256}$$Q_{45,1440}$$Q_{90,\mathsf{\kappa}_{9}}$$Q_{45,7680}$$Q_{45,1536}$$Q_{45,2304}$$Q_{45,11520}$$Q_{45,161280}$$Q_{45,\mathsf{A}_{9}^{5}}$$Q_{136,\mathsf{\Lambda}_{9}}$$Q_{45,\mathsf{A}_{9}^{2}}$
@@
We consider all forms that were not found yet after incrementally exploring all forms until a kissing number of $2\cdot 60$. Concurrent to this work, such a computation was also performed by Gregory Minton [Minton2025] up to a kissing number of $2\cdot 70$, during which he found 2 237 251 031 perfect forms.
There are only four perfect forms with a large kissing number that lead to the remaining nine perfect forms, namely forms that we denote by $Q_{72,D_{9}},Q_{74,256},Q_{90,\mathsf{\kappa}_{9}}$ and $Q_{136,\mathsf{\Lambda}_{9}}$ with a half-kissing number of respectively $72,74,90$ and $136$. The first one $Q_{72,D_{9}}$ and its singularly-connected neighbour $Q_{45,A_{9}}$ correspond to the $D_{9}$ and $A_{9}$ root lattices respectively. The latter two correspond to the lattices $\mathsf{\kappa}_{9}$ [plesken1993constructing] and the laminated lattice $\mathsf{\Lambda}_{9}$ [chaundy1946arithmetic, conway1982laminated] respectively, both sections of the Leech lattice $\Lambda_{24}$ [leech1967notes]. For our exploration the root lattice $A_{9}$ was already part of the starting set, so technically we did not find it at a later stage.
The new perfect forms and their connections in Voronoi’s graph are presented in fig. 4. Being only connected via these high-incidence perfect forms already makes them interesting, but in fact quite some of these lattices have other special properties. All these new perfect forms have the minimal possible kissing number of $2\cdot 45$, but all have a large automorphism group which is typically not the case for such low-incidence forms. The perfect form $Q_{45,\mathsf{A}_{9}^{2}}$ is the unique perfect form that is only connected to the perfect lattice $\mathsf{\Lambda}_{9}$, and corresponds to Coxeter’s lattice $\mathsf{A}_{9}^{2}$ [coxeter1951extreme]. We will see later that this lattice is the only $9$-dimensional perfect lattice whose minimal vectors do not span the lattice. Then there is the perfect form $Q_{45,\mathsf{A}_{9}^{5}}$ [coxeter1951extreme, martinet2004generalization], dual to $Q_{45,\mathsf{A}_{9}^{2}}$, which is only connected in the Voronoi graph via two and three steps with $Q_{90,\mathsf{\kappa}_{9}}$ and $Q_{136,\mathsf{\Lambda}_{9}}$ respectively. We will show later that this lattice is in fact the only so-called hollow perfect lattice in dimension $9$. The perfect forms $Q_{45,\mathsf{A}_{9}^{5}},Q_{45,\mathsf{A}_{9}^{2}}$ and $Q_{45,7680}$ all satisfy $\gamma^{\prime}(Q)^{2}=\lambda_{1}(Q)\cdot\lambda_{1}(Q^{-1})=\frac{16}{5}$, the maximum known value for the Bergé-Martinet invariant $\gamma^{\prime}(Q)$ in dimension $9$ [berge1989probleme]. Along with the perfect form $Q_{45,1440}$, these $4$ forms are both extreme and dual-extreme, i.e., they achieve a local-maximum for the Bergé-Martinet invariant. Furthermore, the perfect form $Q_{45,161280}$ achieves a rather high value of $\gamma^{\prime}(Q_{45,161280})=37/12$, but is not locally-optimal, due to its better neighbour $Q_{45,\mathsf{A}_{9}^{5}}$.
@@ 7.4. Possible kissing numbers in dimension 99
By definition $Q$ lies on the face of the Rhyskov polyhedra $\mathcal{P}_{\lambda_{1}(Q)}^{9}$ defined by the facets corresponding to $\operatorname{Min}(Q)/\{\pm\}$. The incidence of the face is $\frac{1}{2}|\operatorname{Min}(Q)|$ and thus corresponds to half the kissing number of $Q$. This face is either a vertex, i.e., $Q$ is a perfect form, or the face contains a vertex corresponding to a perfect form $Q^{\prime}$. In the first case we know by our full classification that $|\operatorname{Min}(Q)|\in\{45,\ldots,82,84,90,91,99,129,136\}$. Furthermore, we know that $1,\ldots,44$ are attained in dimension $8$ [PerfectDim8] which by a direct-sum construction can be lifted to dimension $9$. In the second case the face can be identified with a non-trivial face of the tangent cone $\mathcal{P}(Q^{\prime})$ at $Q^{\prime}$.
@@
In [vanWoerdenMasterThesis] it was shown that $\operatorname{Vor}(Q_{129})$ can be decomposed into the Voronoi domain $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ of the $8$-dimensional root lattice $\mathsf{E}_{8}$ with $120$ extreme rays and a linearly independent simplicial cone with $9$ extreme rays. As a result, all faces of $\operatorname{Vor}(Q_{129})$ can be described as a sum of a face of $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ and a face of the simplicial cone. If we consider all faces containing $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ we thus obtain non-trivial faces of incidence $120+k$ for $k=0,\ldots,8$. The largest non-trivial face of $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ has incidence $75$, therefore all other faces of $\operatorname{Vor}(Q_{129})$ not containing $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ have incidence at most $75+9=84$.